Can you solve for the following scenario. Thank you!
s Rachels 30 birthday. Five years ago, Rac hel opened a brokerage account when her grandmsother gave her $25,000 for her 25 birthday. Rachel added $2,000 to this account n her 26 birthday, $3,000 on her 27e birthday, $4,000 on her 28 birthday, and $5,000 on her 29h birthday. Rachels goal is to have $400,000 in the account by her 40 birthday ing today, she plans to contribute a fixed amount to the account each year on her birthday. She will make 11 contributions, the first one will occur today, and the final contrib occur on her 40 birthday. Complicating things somewhat is the fact that Rachel plans to withdraw $20,000 from the account on her 35 birthday to finance the down payment home. How large does each o these contrib ut ons have to be r Rachel to reach her goal? Assume the accon t has earned and will continue to earn an effective return of 12% a 国画 IT
Answer #1
Future Value(FV) of Cash Flow on40th Birthday (at end of 15th Year):
FV=(Cash Flow)*((1+i)^(15-N))
i=Interest Rate=12%=0.12
N=Year of Cash Flow
Assume 25th Birthday as 0 year
CASH FLOW IS GIVEN BELOW
N A B=A*(1.12^(15-N) C D=B*(1.12^(15-N)
Year Cash flow FV of Cash Flow Annual Savings Future Value of Savings
25th birthday 0 $25,000 $136,839
26th birthday 1 $2,000 $9,774
27th birthday 2 $3,000 $13,090
28th birthday 3 $4,000 $15,584
29th birthday 4 $5,000 $17,393
TODAY 30th birthday 5 $0 $11,743.93 $36,475
31st birthday 6 $0 $11,743.93 $32,567
32nd birthday 7 $0 $11,743.93 $29,078
33rd birthday 8 $0 $11,743.93 $25,962
34th birthday 9 $0 $11,743.93 $23,180
35th birthday 10 ($20,000) ($35,247) $11,743.93 $20,697
36th birthday 11 $0 $11,743.93 $18,479
37th birthday 12 $0 $11,743.93 $16,499
38th birthday 13 $0 $11,743.93 $14,732
39th birthday 14 $0 $11,743.93 $13,153
40th birthday 15 $0 $11,743.93 $11,744
SUM $157,434 SUM $242,566
TOTAL FV $400,000 (157434+242566)
SUM of Future Values $157,434
Required amount (Goal) $400,000
Additional amount required $242,566 (400000-157434)
Amount of each of 11 annual contributions required beginning today $11,743.93 (Using PMT Function of excel with Rate=12%, Nper=11, Fv=-242566)